Languages for Describing Sheaves of Modules

Alexandra Ralph

For a locally finitely presented category there is a natural language based on a generating set of finitely presented objects. This can be applied to certain categories of sheaves of modules, and is of particular interest when sheaves of modules over sheaves of locally definable scalars are considered.

Address: Department of Mathematics, University of Manchester, Oxford Road, Manchester, M13 9PL, England

E-mail: mbbxfar2@ma.man.ac.uk